OpenStudy (anonymous):

Evaluate \[\sum_{n=2}^{\infty}{\frac{1}{n(n+2)}}\]

OpenStudy (anonymous):

1/n(n+2) = A/n + B/(n+2) A(n+2) + Bn = 1 An + 2A +Bn = 1 A = 0.5 , B = -0.5 sum (1/2n - 1/2(n+2)) from n =2 now observe : sum(1/2n) (from n =2) = 1/4 + 1/6 + 1/8 + 1/10 + 1/12 ... sum(1/2(n+2)) (from n =2) = 1/8 + 1/10 + 1/12.... subtract the second from the first and get 1/4 + 1/6 = 5/12.

OpenStudy (anonymous):

you haven't considered the last terms of both series.

OpenStudy (anonymous):

what do you mean ? i found a formula for this series too now : http://upload.wikimedia.org/wikipedia/en/math/a/b/1/ab110a78934c413b000447358976b4ed.png just subtract the first term (1/3) from this sum

OpenStudy (anonymous):

hmm the integral converges... so I guess the sum also converges.

OpenStudy (anonymous):

yes its 5/12 thanks